# Improving Composed Image Retrieval via Contrastive Learning with Scaling Positives and Negatives

### v1

@@ 4.1.1 Baselines
4.1.2 Training Protocol
We directly use checkpoints released by baseline works as the first stage models to avoid retraining. For the second stage, we calculate image representations for all images before training and only finetune the query encoder using scaled positives and negatives for 5 epochs. For the main results in section 4.2, we only use images in $\Omega$, so it can be fairly compared to any model that uses the original dataset.
Table 1: Average token length calculated by LLAVA Tokenizer Liu et al. (2023a) of the modified text and triplet count statistics for the annotated and generated training sets.
@@ 4.1.2 Training Protocol
FashionIQ Wu et al. (2021) 7.8 18k 16.5 96k
CIRR Liu et al. (2021) 15.4 28k 20.9 128k
4.1.3 Evaluation Datasets
We evaluate our model on two commonly used CIR datasets: FashionIQ Wu et al. (2021) and CIRR Liu et al. (2021).
FashionIQ Wu et al. (2021) consists of 30,134 examples extracted from 77,684 images crawled from fashion websites. These images are categorized into Dress, Shirt, and Top&Tee. The modified text is manually annotated for each pair of reference and target images. As in Baldrati et al. (2022); Zhang et al. (2022); Zhao et al. (2022), we use 18,000 examples for training and 6,016 validation examples for testing since the "real" test set is unavailable.
CIRR Liu et al. (2021) (Composed Image Retrieval on Real-life images) contains 21,552 real-life images from the web taken from $NLVR^{2}$, a popular natural language reasoning dataset. CIRR contains 36,554 examples, of which 28,225 examples are used for training, 4,181 for validation, and 4,148 for testing. In addition, the images in this dataset are divided into several semantically similar groups to evaluate $R_{\rm subset}$@$K$ metric (see below).
Table 2: Evaluation results of various models on FashionIQ. The best results are in boldface.
@@ 4.1.3 Evaluation Datasets
CIRPLANT Liu et al. (2021) w/o VLP 17.45 40.41 17.53 38.81 21.64 45.38 18.87 41.53 30.20
ARTEMIS Delmas et al. (2022) w/o VLP 27.16 52.40 21.78 43.64 29.20 54.83 26.05 50.29 38.17
ComqueryFormer Xu et al. (2023b) w/o VLP 28.85 55.38 25.64 50.22 33.61 60.48 29.37 55.36 42.37
PL4CIR Zhao et al. (2022) CLIP 33.60 58.90 39.45 61.78 43.96 68.33 39.02 63.00 51.01
TG-CIR Wen et al. (2023) CLIP 35.55 59.44 40.24 62.37 43.65 67.36 39.81 63.06 51.44
@@
CLIP4CIR Baldrati et al. (2022) CLIP 38.18 62.67 44.01 64.57 45.39 69.56 42.52 65.60 54.06
@@
Table 3: Performance comparison of various models on CIRR. The best results are in boldface.
@@
CIRPLANT Liu et al. (2021) w/o VLP 19.55 52.55 68.39 92.38 39.20 63.03 79.49 45.88
ARTEMIS Delmas et al. (2022) w/o VLP 16.96 46.10 61.31 87.73 39.99 62.20 75.67 43.05
ComqueryFormer Xu et al. (2023b) w/o VLP 25.76 61.76 75.90 95.13 51.86 76.26 89.25 56.81
TG-CIR Wen et al. (2023) CLIP 45.23 78.34 87.13 97.30 72.84 89.25 95.13 75.59
@@
CLIP4CIR Baldrati et al. (2022) CLIP 42.80 75.88 86.26 97.64 70.00 87.45 94.99 72.94
@@
4.1.4 Evaluation Metrics
$Recall$@$K$ ($R$@$K$) is the proportion of queries for which the retrieved top K images include the correct target image. $Recall_{\rm subset}$@$K$ ($R_{\rm subset}$@$K$) is nearly the same as $R$@$K$ but the model only retrieves inside the semantically similar group of the reference image. For the FashionIQ dataset, following previous works Bai et al. (2023); Liu et al. (2023c); Baldrati et al. (2022), we evaluate our model through $R$@$K$ ($K=10,50$) on the original protocol. As in  Zhao et al. (2022), we also report the mean of all $R$@$K$ scores as Rmean. For the CIRR dataset, following previous works Baldrati et al. (2022); Liu et al. (2023b), we evaluate our model through $R$@$K$ ($K=1,5,10,50$) and $R_{\rm subset}$@$K$ ($K=1,2,3$). As in  Bai et al. (2023), we also report ($R$@$5$+$R_{\rm subset}$@$1$)/2 as Rmean.
4.1.5 Implementation Details
@@
We analyze the modified text in the two datasets using LLAVA tokenizer Liu et al. (2023a) and count the average annotated token length in Table 1. For FashionIQ, we set $type$ to the name of the split, i.e., dress/shirt/top tee, and $k$ to 5. For CIRR and Conceptual Caption, we set $type$ to "image" and $k$ to 10 in the image captioning template. The detailed data statistics for both the generated and annotated triplets are provided in Table 1. For both datasets, we set $c_{0}$ to 10000. We set $c_{1}$ to 20000 for FashionIQ and 15000 For CIRR.
@@
Figure 4: Discussion of the core components in the method. The results shown in the figures are on the validation set.
4.2 Main Results
We compare our method against the following baseline methods: CIRPLANT Liu et al. (2021), ARTEMIS Delmas et al. (2022), ComqueryFormer Xu et al. (2023b), PL4CIR Zhao et al. (2022), TG-CIR Wen et al. (2023), CLIP4CIR Baldrati et al. (2022), BLIP4CIR Liu et al. (2023c), SPRC Bai et al. (2023). Details about these models can be found in supplementary materials. We abbreviate the method of scaling positive examples as SP, the method of scaling negative examples as SN, and the superposition of the two methods as SPN.
@@ 4.2 Main Results
Table 3 illustrates the comparison between our model and other recent studies on CIRR. It shows that SPN also improves the performance of all four baseline models. SPN increases the R@1 of TG-CIR by 4.5%, CLIP4CIR by 5.9%, BLIP4CIR by 3.7%, and SPRC by 6%. SPN improves the R@5 of TG-CIR by 1%, CLIP4CIR by 2.9%, BLIP4CIR by 1.4%, and SPRC by 2.1%. This proves our method works well for images in both general and fashion scenes. SPN promotes the Rsubset@1 of TG-CIR by 3.5%, CLIP4CIR by 5.6%, BLIP4CIR by 1%, and SPRC by 1.1%. The objective of the subset test is to justify whether the model can distinguish between hard negative examples Liu et al. (2021). Such a boost indicates that our model can learn more fine-grained representations than the base model, thus distinguishing harder negative examples.
Table 4: Ablation results on CLIP4CIR.
@@ 4.3 Ablation Study
The MLLM we use can be replaced with any model that can generate captions for images, so we try three representative models LLAVA Liu et al. (2023a), BLIP Li et al. (2022), and BLIP-2 Li et al. (2023) in Table 4(b). We find that LLAVA, with great instruction fine-tuning, works best among the three models. But surprisingly, BLIP works better than BLIP-2. This suggests that BLIP-2’s ability to follow image captioning instructions is not very good. At the same time, using different MLLMs consistently yields better results than w/o SP, indicating that our method is insensitive to different MLLMs.
@@
SP allows for constructing many triplets based on images, so we consider exactly how many additional triplets on top of the existing ones work best. As shown in Fig 4(c), as the number of positive examples rises, the effect of the model increases and then decreases, with the best results when increasing nearly 60% of the number of original triplets, that is 12k for FashionIQ and 16k for CIRR.
@@
In Fig.4(d), we explore four methods for constructing image pairs. The first method involves selecting target images with the highest similarity to the reference image. The second method entails choosing target images with moderate similarity to the reference image. The third method focuses on selecting target images with the lowest similarity to the reference image, while the fourth method involves selecting target images randomly from the entire set. Our findings indicate that the second method consistently produces superior results across both datasets. This observation can be attributed to the higher quality of triplets generated through this selection approach compared to the others.
@@
Table 5: Zero-shot results.
@@
CLIP Radford et al. (2021) 19.04 35.03 12.65 38.41 34.29
PIC2WORD Saito et al. (2023) 24.70 43.70 23.90 51.90 -
SEARLE-OTI Baldrati et al. (2023) 27.61 47.90 24.87 52.31 53.80
@@
4.4 Results of Zero-Shot CIR
@@
Neither of these baselines uses an in-domain image dataset for training. Therefor, we also utilize images from the out-of-domain dataset Conceptual Caption (CC3M), comprised of 3.3 million image-caption pairs from the Internet, to generate positive examples for a fair comparison with PIC2WORD. Specifically, we randomly select the 50k images in CC3M to construct the CIR dataset due to computing resource limitations. The number of 50k images is equal to 1.7% of that PIC2WORD used and 50% of that SEARLE used. We abbreviate our model trained with this setting as SPN-CC. As shown in Table 5, SPN-CC gets the best results while using the least amount of images. This suggests that, given a collection of images out of the domain, our method can automatically construct appropriate triplets and train an acceptable model in the zero-shot setting. We also explore the setting of in-domain images, i.e., those in FashionIQ and CIRR, and abbreviate the model trained with this setting as SPN-IN. Under this setting, SPN-IN yields better results than SPN-CC using out-of-domain data. This suggests that if in the future we need to do a composed image retrieval task for a new scene but with few labeling costs, an accepted solution is to use our SPN method to automatically construct the positive examples within this scene and train a model from scratch.
Figure 5: Comparison of retrieval results between the CLIP4CIR model w/o and w SPN.
4.5 Case Study
@@
Fig.5 indicates the retrieval cases of CLIP4CIR w/o and w SPN. The first example is selected from CIRR, and the second one is from FashionIQ. For both examples, we can find that using SPN allows us to learn more of the rarer concepts (e.g. "llama", "logo"), thus enhancing the base model. In the meantime, we can find that the base model has difficulty in retrieving the correct image when the reference and target images are very different (e.g., "panda" and "llama", "dress" and "tee"), and SPN narrows this gap. More examples can be found in supplementary materials.
5 Conclusion
@@
The Composed Image Retrieval (CIR) task uses a composed query to retrieve target images. While existing methods have achieved impressive results, limited labeled data and contrastive learning with in-batch negative sampling limit the performance of their methods. To address these problems, we first propose a data generation method using a multi-modal large language model to scale positives. We then propose a two-stage fine-tuning framework to scale negatives, introducing static representations of negatives in the second stage. These improvements are plug-and-play, enhancing existing CIR models without architecture changes. Extensive experiments show that we obtain state-of-the-art results on the FashionIQ and CIRR datasets. Moreover, our method could be applied to zero-shot composed image retrieval, offering a novel solution for unannotated CIR scenarios.

### v2

@@ 4.1.1. Baselines
4.1.2. Training Protocol
We directly use the checkpoints released by the baseline as first stage models to avoid retraining. For the second stage, we calculate image representations for all images before training and only finetune the query encoder using scaled positives and negatives for 5 epochs. For the main results in section 4.2, we only use images in $\Omega$, so it can be fairly compared to any model that uses the original dataset.
Table 1. Average token length calculated by LLAVA Tokenizer (Liu et al., 2023a) of the modified text and triplet count statistics for the annotated and generated training sets.
@@ 4.1.2. Training Protocol
FashionIQ (Wu et al., 2021) 7.8 18k 16.5 96k
CIRR (Liu et al., 2021) 15.4 28k 20.9 128k
4.1.3. Evaluation Datasets
We evaluate our model on two commonly used CIR datasets, fashion-domain FashionIQ (Wu et al., 2021) and open-domain CIRR (Liu et al., 2021).
Table 2. Evaluation results of various models on FashionIQ. The best results are in boldface.
@@ 4.1.3. Evaluation Datasets
CIRPLANT (Liu et al., 2021) w/o VLP 17.45 40.41 17.53 38.81 21.64 45.38 18.87 41.53 30.20
ARTEMIS (Delmas et al., 2022) w/o VLP 27.16 52.40 21.78 43.64 29.20 54.83 26.05 50.29 38.17
ComqueryFormer (Xu et al., 2023b) w/o VLP 28.85 55.38 25.64 50.22 33.61 60.48 29.37 55.36 42.37
PL4CIR (Zhao et al., 2022) CLIP 33.60 58.90 39.45 61.78 43.96 68.33 39.02 63.00 51.01
TG-CIR (Wen et al., 2023) CLIP 35.55 59.44 40.24 62.37 43.65 67.36 39.81 63.06 51.44
@@
CLIP4CIR (Baldrati et al., 2022) CLIP 38.18 62.67 44.01 64.57 45.39 69.56 42.52 65.60 54.06
@@
Table 3. Performance comparison of various models on CIRR. The best results are in boldface.
@@
CIRPLANT (Liu et al., 2021) w/o VLP 19.55 52.55 68.39 92.38 39.20 63.03 79.49 45.88
ARTEMIS (Delmas et al., 2022) w/o VLP 16.96 46.10 61.31 87.73 39.99 62.20 75.67 43.05
ComqueryFormer (Xu et al., 2023b) w/o VLP 25.76 61.76 75.90 95.13 51.86 76.26 89.25 56.81
TG-CIR (Wen et al., 2023) CLIP 45.23 78.34 87.13 97.30 72.84 89.25 95.13 75.59
@@
CLIP4CIR (Baldrati et al., 2022) CLIP 42.80 75.88 86.26 97.64 70.00 87.45 94.99 72.94
@@
4.1.4. Evaluation Metrics
$Recall$@$K$ ($R$@$K$) is the proportion of queries for which the retrieved top K images include the correct target image. $Recall_{\rm subset}$@$K$ ($R_{\rm subset}$@$K$) is nearly the same as $R$@$K$ but the model only retrieves inside the semantically similar group of the reference image. For the FashionIQ dataset, following previous works (Bai et al., 2023; Liu et al., 2023c; Baldrati et al., 2022), we evaluate our model through $R$@$K$ ($K=10,50$) on the original protocol. As in  (Zhao et al., 2022), we also report the mean of all $R$@$K$ scores as Rmean. For the CIRR dataset, following previous works (Baldrati et al., 2022; Liu et al., 2023b), we evaluate our model through $R$@$K$ ($K=1,5,10,50$) and $R_{\rm subset}$@$K$ ($K=1,2,3$). As in  (Bai et al., 2023), we also report ($R$@$5$+$R_{\rm subset}$@$1$)/2 as Rmean.
4.1.5. Implementation Details
@@
We analyze the modified texts in the two datasets using the LLaVA tokenizer (Liu et al., 2023a) and count the average length of the annotated token in Table 1. For FashionIQ, we set $type$ to the name of the split, i.e., dress/shirt/top tee, and $k$ to 5. For CIRR and Conceptual Caption, we set $type$ to ”image” and $k$ to 10 in the image captioning template. The detailed data statistics for both the generated and annotated triplets are provided in Table 1. For both datasets, we set $c_{0}$ to 10000. We set $c_{1}$ to 20000 for FashionIQ and 15000 for CIRR.
@@
Figure 4. Discussion of the core components in the method. The results shown in the figures are on the validation set.
[…]
[Discussion of the core components in the method.]Discussion of the core components in the method.
4.2. Main Results
We compare our method with the following baseline methods: CIRPLANT (Liu et al., 2021), ARTEMIS (Delmas et al., 2022), ComqueryFormer (Xu et al., 2023b), PL4CIR (Zhao et al., 2022), TG-CIR (Wen et al., 2023), CLIP4CIR (Baldrati et al., 2022), BLIP4CIR (Liu et al., 2023c), SPRC (Bai et al., 2023). Details about these models can be found in the supplementary material. We abbreviate the method of scaling positive examples as SP, the method of scaling negative examples as SN, and the superposition of the two methods as SPN.
@@ 4.2. Main Results
Table 3 illustrates the comparison between our model and other recent studies on CIRR. It shows that SPN also improves the performance of all four baseline models. SPN increases the R@1 of TG-CIR by 4.5%, CLIP4CIR by 5.9%, BLIP4CIR by 3.7%, and SPRC by 6%. SPN improves R@5 of TG-CIR by 1%, CLIP4CIR by 2.9%, BLIP4CIR by 1.4%, and SPRC by 2.1%. This shows that our method works well for images in both general and fashion scenes. SPN promotes the Rsubset@1 of TG-CIR by 3.5%, CLIP4CIR by 5.6%, BLIP4CIR by 1%, and SPRC by 1.1%. The objective of the subset test is to justify whether the model can distinguish between hard negative examples (Liu et al., 2021). Such a boost indicates that our model can learn more fine-grained representations than the base model, thus distinguishing harder negative examples.
Table 4. Ablation results on CLIP4CIR.
@@ 4.3. Ablation Study
The MLLM we use can be replaced with any model that can generate captions for images, so we try three representative models LLaVA (Liu et al., 2023a), BLIP (Li et al., 2022), and BLIP-2 (Li et al., 2023) in Table 4(b). We find that LLaVA, with great instruction fine-tuning, works best among the three models. But surprisingly, BLIP works better than BLIP-2. This suggests that BLIP-2’s ability to follow image captioning instructions is not very well. At the same time, using different MLLMs consistently yields better results than w/o SP, indicating that our method is insensitive to different MLLMs.
@@
SP allows for constructing plenty of triplets based on images, so we consider exactly how many additional triplets on top of the existing ones work best. As shown in Fig 4(c), as the number of positive examples rises, the effect of the model increases and then decreases, with the best results when increasing nearly 60% of the number of original triplets, that is 12k for FashionIQ and 16k for CIRR.
@@
In Fig.4(d), we explore four methods for constructing image pairs. The first method involves selecting target images with the highest similarity to the reference image. The second method entails choosing target images with moderate similarity to the reference image. The third method focuses on selecting target images with the lowest similarity to the reference image, while the fourth method involves selecting target images randomly from the entire set. Our findings indicate that the second method consistently produces superior results across both datasets. This can be attributed to the fact that this approach is closest to the way triplets are constructed in the original datasets.
@@
Table 5. Zero-shot results.
@@
CLIP (Radford et al., 2021) 19.04 35.03 12.65 38.41 34.29
PIC2WORD (Saito et al., 2023) 24.70 43.70 23.90 51.90 -
SEARLE-OTI (Baldrati et al., 2023) 27.61 47.90 24.87 52.31 53.80
@@
4.4. Results of Zero-Shot CIR
@@
Neither of these baselines uses an in-domain image dataset for training. Therefore, we also utilize images from the out-of-domain dataset Conceptual Caption (CC3M), consisting of 3.3 million image-caption pairs from the Internet, to generate positive examples for a fair comparison with PIC2WORD. Specifically, we randomly select the 50k images in CC3M to construct the CIR dataset. The number of 50k images is equal to 1.7% of that PIC2WORD use and 50% of that SEARLE use. We abbreviate our model trained with this setting as SPN-CC. As shown in Table 5, SPN-CC gets the best results while using the least amount of images. This suggests that, given a collection of images out of the domain, our method can automatically construct appropriate triplets and train an acceptable model in the zero-shot setting. We also explore the setting of in-domain images, i.e., those in FashionIQ and CIRR, and abbreviate the model trained with this setting as SPN-IN. Under this setting, SPN-IN yields better results than SPN-CC using out-of-domain data. This suggests that if we need to do a CIR task for a new scene with few labeling costs, SPN can automatically construct the positive examples within this scene and train a model from scratch.
4.5. Case Study
@@
Fig.5 indicates the retrieval cases of CLIP4CIR w/o and w SPN. The first example is from CIRR, and the second one is from FashionIQ. For both examples, we find that using SPN allows us to learn more of the rarer concepts (e.g. ”llama”, ”logo”), thus enhancing the base model. In the meantime, we can find that the base model has difficulty in retrieving the correct image when the reference and target images are very different (e.g., ”panda” and ”llama”, ”dress”, and ”tee”), and SPN narrows this gap. More examples can be found in the supplementary material.
Figure 5. Comparison of retrieval results between the CLIP4CIR model w/o and w SPN.
[…]
[Comparison of retrieval results between the CLIP4CIR model w and w/o SPN.]Comparison of retrieval results between the CLIP4CIR model w and w/o SPN.
5. Conclusion
@@
The Composed Image Retrieval (CIR) task uses a composed query to retrieve target images. While existing methods have achieved impressive results, limited labeled data and contrastive learning with in-batch negative sampling hinder the performance of these methods. We propose a data generation method using a multi-modal large language model to populate positives, followed by a two-stage fine-tuning framework for scaling negatives, introducing static negative representations in the second stage. These improvements integrate seamlessly into existing CIR models. Extensive experiments demonstrate state-of-the-art results on the FashionIQ and CIRR datasets, and our approach can also be applied to the zero-shot composed image retrieval, providing a novel solution for unannotated CIR scenarios.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (No. U23B2056), in part by the Fundamental Research Funds for the Central Universities, and in part by the State Key Laboratory of Complex & Critical Software Environment.